# Gamma-ray Bursts as Distance Indicators by a Machine Learning Approach

### v1

@@ Abstract
Gamma-ray bursts (GRBs) can be probes of the early universe, but currently, only 26% of GRBs observed by the Neil Gehrels Swift Observatory GRBs have known redshifts ($z$) due to observational limitations. To address this, we estimated the GRB redshift (distance) via a supervised machine learning model that uses optical afterglow observed by Swift and ground-based telescopes. The inferred redshifts are strongly correlated (a Pearson coefficient of 0.93) with the observed redshifts, thus proving the reliability of this method. The inferred and observed redshifts allow us to estimate the number of GRBs occurring at a given redshift (GRB rate) to be 7.6-8 $yr^{-1}Gpc^{-1}$ for $2.31.9<z<$ 2.3. Since GRBs come from the collapse of massive stars, we compared this rate with the star formation rate highlighting a discrepancy of a factor of 3 at $z<1$.
@@ 1 Introduction
These broad distributions hinder the search for an effective strategy to derive redshifts reliably from the observational features. It is essential to determine how these distributions vary with the redshift, namely, the functional form or their cosmological evolution (Dainotti et al., 2013b; Petrosian et al., 2015; Dainotti et al., 2015a), before attacking any of the above problems. To determine these functional forms, we need large samples with known redshifts and well-defined observational selection criteria. This allows for a precise determination of cosmological evolutions.
@@
Interestingly, Petrosian et al. (2015); Yu et al. (2015); Pescalli et al. (2016) and Tsvetkova et al. (2017) discovered a difference between the evolution of the LGRB and the global SFR at low redshifts ($10<z<$ 1). However, Pescalli et al. (2016) found smaller differences between GRB and star FR. We focus on this ongoing debate and determine the true GRB rate. Our goal with the machine learning (ML) methodology is to increase the number of LGRBs with inferred redshift considerably (we will almost double the sample of Swift GRBs with $z$ if we consider both X-rays and optical data) so that we can solve the above debate.
@@
As mentioned above, using correlations has allowed us to obtain pseudo-redshifts for larger samples for cosmological studies (Lloyd-Ronning, 2022; Atteia, 2003; Yonetoku et al., 2004; Kocevski & Liang, 2006; Tsutsui et al., 2011; Dainotti et al., 2011; Zhang & Wang, 2018), but the inferred redshifts are not yet accurate. We propose a very effective method using ML tools for determining redshifts.
In §2, we detail our GRB sample used for the redshift inference procedure via ML methods, described in §A. In §3, we describe how the chosen procedures reproduce the observed spectroscopic redshifts. In §3.2, we compare the results from applying ML methods to both predicted redshift ($z_{pred}$) and observed redshift ($z_{obs}$) to obtain the LF and density rate evolution. We also compare our density rate evolution with the established results in the literature. A summary and discussion of the impact on future work are given in §4.
Figure 1: The scatter matrix plot of the features at play, some showing strong correlations, while others show weak or no correlation. Each variable (the prompt $Fluence$, $T_{90}$, the prompt $Peakflux$, the plateau optical flux ($\log(F_{a})$) its duration ($\log(T_{a})$) the hydrogen column density ($\log(NH)$), the prompt $Photon$ $Index$, the temporal and spectral indices, $\alpha$, and $\beta$, of the optical afterglow and plateau emissions, respectively, is plotted against every other. The stars next to the correlation value denote the significance level. One, two, or three stars denote the corresponding variables’ correlation is significant at 10%, 5% and 1% levels, respectively.
@@ 2 The sample
In about 42% of cases, the afterglow presents the plateau emission, a flat part of the GRB lightcurves, whose parameters we include as explanatory variables or predictors of their redshifts.
The data sets are composed of 179 GRBs with optical plateaus from the Swift Ultraviolet/Optical Telescope (UVOT) and from 455 ground-based telescopes/detectors, e.g., the Subaru Telescope, Gamma-ray Burst Optical/Near-IR Detector (GROND), Reionization and Transients InfraRed camera/telescope (RATIR), the MITSuME (Kotani et al., 2005), observing from 1997 May to 2021 May. Fig. 1 shows a scatter matrix plot of all the variables (including the redshift). This matrix includes the logarithms of the redshift, prompt $Fluence$, $T_{90}$, the prompt $Peakflux$, the peak flux computed in 2 second, the time at the end of the plateau emission $T_{a}$, its correspondent flux, $F_{a}$, and the column density, $NH$, while the following variables in linear scale: the prompt $Photon$ $Index$, and the spectral index (assuming a simple power-law spectrum), $\beta$, of the plateau emission, and the temporal decay index after the plateau, $\alpha$. As it is clear from the Fig. 1, some of these scatter diagrams show clear linear phenomenological relations such as the bi-dimensional (Dainotti et al., 2008) relation in optical between the plateau luminosity and rest frame duration, $L_{a}$ and $T^{*}_{a}$, and its extension in three dimensions by adding the prompt peak luminosity, $L_{peak}$ (Dainotti et al., 2016; Dainotti & Del Vecchio, 2017a; Dainotti et al., 2020a), and the various spectral hardness-luminosity relations (Collazzi & Schaefer, 2008). For recent reviews on correlations see Dainotti & Del Vecchio (2017b); Dainotti et al. (2018); Dainotti & Amati (2018). In addition to these relations, we also note a strong relation between prompt $Fluence$ and $Peakflux$.
@@ 2.1 The sample data cut and the MICE algorithm
Figure 2: The missing data in our sample. The pink boxes show the missing GRBs in relation to a given variable presented in the upper x-axis, while the blue boxes indicate GRBs with no missing data for a given GRB variable presented in the top axis. The bottom axis shows the number of missing GRBs per variable. The left axis represents the number of observations that have missing data for a specific set of features. The right axis represents the number of features that are missing for that row. For example, starting from the first row, there are 126 GRBs with no missing data, 14 GRBs with missing data in $\log(Peakflux)$ data, 2 GRBs with missing data in $\log(NH)$, and $\log(Peakflux)$ error, and so on.
@@
MICE can impute missing values for multiple variables using the data set with complete variables. However, the missing data has to be assumed to be missing at random (MAR) (Rubin, 1976; Schafer & Graham, 2002). Here, we use the method of predictive mean matching (PMM)(Little & Rubin, 2019), named ‘midastouch’. In PMM, missing features are initialized with their mean values, and then their estimates are obtained by training a model on the rest of the complete data. To reduce the imputation’s randomness, we will impute the missing variables 20 times for each MICE iteration instead of the usual practice of 10 times (Van Buuren & Groothuis-Oudshoorn, 2011). We tested the imputed and original distributions with the Kolmogorov-Smirnov test, which yielded a significance level of 0.5. Previously, Gibson et al. (2022) applied MICE to impute missing values of Active Galactic Nuclei. There, we demonstrated that this imputation method does not add any additional uncertainty to the variable’s original distribution and helps strengthen the final training set used for the ML application by enlarging it.
@@ 3 Results
3.1 The Machine learning results
We present the results derived with the ensemble method, the SuperLearner, which has the advantage of combining several ML methods into a single model and leveraging the predictive power of each model. This is the first time SuperLearner has been applied to GRBs. SuperLearner computes coefficients for each learner based on their performance within the ensemble. Each model and the Superlearner are discussed in §A.1. In the left panel of Fig. 3, we show the observed redshift ($z_{obs}$) versus the predicted redshift ($z_{pred}$) via a 10-fold cross-validation technique averaged over 100 runs (see §A.3). The average mean squared error, MSE, is defined as the $<(z_{pred}-z_{obs})^{2}>$ where the symbol <>absent<>< > indicates the average between $z_{pred}$ vs $z_{obs}$. The obtained MSE is $0.21$, with a high Pearson correlation ($r$) of $0.93$ and a very small average bias, defined as the mean $<(z_{pred}-z_{obs})>$ of $0.02$.
@@ 3.1 The Machine learning results
In the right panel of Fig. 3 we have shown the relative influence of the variables obtained with SuperLearner (see §A.3 for details). The relative importance of features is an average of the importance given by a local linear approximation of predictions. The variable $\log(NH)$ obtained the highest influence in predicting the redshift. This variable gives the NH obtained from an automated time-averaged spectral fit of the Swift XRT data (Evans et al., 2009). The high influence of this variable can be understood because the more distant a GRB is, the more neutral hydrogen it will encounter. Thus, the redshift of a GRB depends on the calculated $\log(NH)$ value. The second, third, fourth and fifth most predictive variables are $\log(F_{a})$, $\alpha$, $\log(Peakflux)$, $\log(T_{a})$. The variables related to prompt and the plateau emission, such as the time at the end of the plateau ($log(T_{a})$) and its correspondent flux $\log(F_{a})$, and the prompt $Peakflux$ are among the best predictors, thus recovering the key parameters of the three-dimensional Dainotti correlation (named the GRB fundamental plane relation), the correlation between the peak prompt luminosity, the luminosity at the end of the plateau emission and the rest end time of the plateau emission (Dainotti et al., 2016, 2017a, 2020a, 2020b, 2022).
@@ 3.2 Luminosity function and density rate evolution
[…]
@@
However, the values of these luminosity breaks are lower than those in Pescalli et al. (2016); Dong et al. (2022); Yu et al. (2015); Petrosian et al. (2015); Wanderman & Piran (2015, 2010), where the average log luminosity break is $51$. This is because we use plateau luminosity for deriving the LF. The difference between the flux of the prompt and plateau is of the order of $\sim 10^{5}$. If we account for this in our luminosity break, then our values aligns with the literature.
@@
The density rates (see Fig. 5) are calculated using the same 98 GRBs as in the LF. The plots compare our density rate and those in the literature (upper panel of Fig. 5) and our density rate with SFR (lower panel of Fig. 5). We found the actual rate to be $7.6-8$ Gpc $yr^{-1}$ for $2.31.9<z<$ 2.3 for the $z_{obs}$ and $z_{pred}$, respectively, without an arbitrary normalization. We have renormalized the rate (with a factor of 20) to account for our subset of LGRBs, compared to all LGRBs detected by Fermi and Swift, from the first detection of Fermi (14th July 2008) to 30th May 2023. We have quantitatively shown that the SFR does not match the GRB rate at low redshift ($z<1$). This discrepancy is of a factor of $\sim$3, comparing with Hopkins & Beacom (2006). The SFR that better fit our GRB data is Hopkins & Beacom (2006); Graham & Schady (2016), where the agreement runs from $z=0.8$ until $z=4.8$.
@@
The GRB rate does not follow the SFR at $z<1$, supporting the previous studies with a quantitative estimate. Although for most of the SFR compared here, the GRB rate does not follow the SFR at low-$z$, in the case of Thompson et al. (2006), there is no disagreement due to the very large errorbars. In general, the nature of this discrepancy can be due to selection biases (Petrosian et al., 2015) or to metallicity (Yoon et al., 2006; Ghirlanda & Salvaterra, 2022). The last two papers explain this excess due to the increased metallicity with cosmic time, which prevents a larger fraction of massive stars from producing GRBs. Ghirlanda & Salvaterra (2022) found that a metallicity threshold of $8.612+log(O/H)<=$ 8.6 accounts for the derived GRB efficiency evolution. They verified their results, computing the stellar masses of galaxies with metallicities below the threshold value found and compared them with the observed masses of a sample of GRB hosts claimed to be complete. The maximum and average stellar masses are consistent with the measured masses of the hosts and their redshift evolutions. However, these results depend on the claim of the completeness of the sample of the GRB hosts, which can be debated. Another relevant possibility is the low-$z$ GRBs have the same non-collapsar origin (Petrosian & Dainotti, 2023) of the SGRBs. The resulting LF and density rate evolutions are consistent with the ones in the literature and are also consistent between the ones derived from $z_{obs}$ and $z_{pred}$. We have shown in the right panel of Fig. 4 that adding the estimated redshift improves the derivation of the break of the LF by 90%, namely $46.6\pm 0.3$ vs $47.2\pm 0.03$.
@@
Figure 5: The orange and purple dots show the density rates obtained from $z_{obs}$ and $z_{pred}$, respectively. Top Panel: the density rate evolution for our sample, compared with other published density rate evolutions. Bottom Panel: The density rate evolution for our sample compared with the published SFR.

### v2

# Gamma-ray Bursts as Distance Indicators by a Statistical Learning Approach

@@ Abstract
[…]
@@ 1 Introduction
These broad distributions hinder the search for an effective strategy to derive redshifts reliably from the observational predictors. It is essential to determine how these distributions vary with the redshift, namely, the functional form or their cosmological evolution (Dainotti et al., 2013b; Petrosian et al., 2015; Dainotti et al., 2015a), before attacking any of the above problems. To determine these functional forms, we need large samples with known redshifts and well-defined observational selection criteria. This allows for a precise determination of cosmological evolutions.
@@
Interestingly, Petrosian et al. (2015); Yu et al. (2015); Pescalli et al. (2016) and Tsvetkova et al. (2017) discovered a difference between the evolution of the LGRB and the global SFR at low redshifts ($10<z<$ 1). However, Pescalli et al. (2016) found smaller differences between GRB and star FR. We focus on this ongoing debate and determine the true GRB rate. Our goal with the statistical learning methodology is to increase the number of LGRBs with inferred redshift considerably (we will almost double the sample of Swift GRBs with $z$ if we consider both X-rays and optical data) so that we can solve the above debate.
@@
As mentioned above, using correlations has allowed us to obtain pseudo-redshifts for larger samples for cosmological studies (Aldowma & Razzaque, 2024; Lloyd-Ronning, 2022; Atteia, 2003; Yonetoku et al., 2004; Kocevski & Liang, 2006; Tsutsui et al., 2011; Dainotti et al., 2011; Zhang & Wang, 2018), but the inferred redshifts are not yet accurate. We propose a very effective method using statistical learning tools for determining redshifts.
In §2, we detail our GRB sample used for the redshift inference procedure via methods, described in §A. In §3, we describe how the chosen procedures reproduce the observed spectroscopic redshifts. In §3.2, we compare the results from applying statistical learning methods to both predicted redshift ($z_{pred}$) and observed redshift ($z_{obs}$) to obtain the LF and density rate evolution. We also compare our density rate evolution with the established results in the literature. A summary and discussion of the impact on future work are given in §4.
Figure 1: The scatter matrix plot of the predictors at play, some showing strong correlations, while others show weak or no correlation. Each variable (the prompt $Fluence$, $T_{90}$, the prompt $Peakflux$, the plateau optical flux ($\log(F_{a})$) its duration ($\log(T_{a})$) the hydrogen column density ($\log(NH)$), the prompt $Photon$ $Index$, the temporal and spectral indices, $\alpha$, and $\beta$, of the optical afterglow and plateau emissions, respectively, is plotted against every other. The stars next to the correlation value denote the significance level. One, two, or three stars denote the corresponding variables’ correlation is significant at 10%, 5% and 1% levels, respectively.
@@ 2 The sample
In about 42% of cases, the afterglow presents the plateau emission, a flat part of the GRB lightcurves, whose parameters we include as predictors of their redshifts.
The data sets are composed of 179 GRBs with optical plateaus from the Swift Ultraviolet/Optical Telescope (UVOT) and from 455 ground-based telescopes/detectors, e.g., the Subaru Telescope, Gamma-ray Burst Optical/Near-IR Detector (GROND), Reionization and Transients InfraRed camera/telescope (RATIR), the MITSuME (Kotani et al., 2005), observing from 1997 May to 2021 May. Fig. 1 shows a scatter matrix plot of all the variables (including the redshift). This matrix includes the logarithms of the redshift, prompt $Fluence$, $T_{90}$, the prompt $Peakflux$, the peak flux computed in 2 second, the time at the end of the plateau emission $T_{a}$, its correspondent flux, $F_{a}$, and the column density, $NH$, while the following variables in linear scale: the prompt $Photon$ $Index$, and the spectral index (assuming a simple power-law spectrum), $\beta$, of the plateau emission, and the temporal decay index after the plateau, $\alpha$. As it is clear from the Fig. 1, some of these scatter diagrams show clear linear phenomenological relations such as the bi-dimensional Dainotti et al. (2008) relation in optical (Dainotti et al., 2020b) between the plateau luminosity and rest frame duration, $L_{a}$ and $T^{*}_{a}$, and its extension in three dimensions (3D) by adding the prompt peak luminosity, $L_{peak}$ (Dainotti et al., 2016; Dainotti & Del Vecchio, 2017a; Dainotti et al., 2020a), and the various spectral hardness-luminosity relations (Collazzi & Schaefer, 2008). The bi-dimensional $L_{a}$ and $T^{*}_{a}$ relation stems from the $\log F_{a}-\log T_{a}$ relation between the flux at the end of the plateau emission, $F_{a}$, and its duration, $T_{a}$ in the observer frame, as the predictors shown in Fig.1, has a Pearson correlation of $-0.643$. The 3D correlation can be seen through the observer frame relation between the $\log Peakflux$ vs $\log F_{a}$. Here, the Pearson correlation is small since in the 3D relation (Dainotti et al., 2022) the peak flux is given in energy flux (erg/cm2s) and not as a photon flux (photons/cm2s) as in this work. For these relations, we acknowledge that from previous works (Dainotti et al., 2013b, 2015a, 2017a, 2020b; Levine et al., 2022; Dainotti et al., 2022) we have shown that these relations are intrinsic and not induced by detector threshold truncations. For recent reviews on correlations see Dainotti & Del Vecchio (2017b); Dainotti et al. (2018); Dainotti & Amati (2018). In addition to these relations, we also note a strong relation between prompt $Fluence$ and $Peakflux$.
@@ 2.1 The sample data cut and the MICE algorithm
Figure 2: The missing data in our sample. The pink boxes show the missing GRBs in relation to a given variable presented in the upper x-axis, while the blue boxes indicate GRBs with no missing data for a given GRB variable presented in the top axis. The bottom axis shows the number of missing GRBs per variable. The left axis represents the number of observations that have missing data for a specific set of predictors. The right axis represents the number of predictors that are missing for that row. For example, starting from the first row, there are 126 GRBs with no missing data, 14 GRBs with missing data in $\log(Peakflux)$ data, 2 GRBs with missing data in $\log(NH)$, and $\log(Peakflux)$ error, and so on. The corner number on the right of the X-axis (186) shows the total number of missing predictors in the data set, which is the sum of the numbers shown along the axis
@@
MICE can impute missing values for multiple variables using the data set with complete variables. However, the missing data has to be assumed to be missing at random (MAR) (Rubin, 1976; Schafer & Graham, 2002). Here, we use the method of predictive mean matching (PMM)(Little & Rubin, 2019), named ‘midastouch’. In PMM, missing predictors are initialized with their mean values, and then their estimates are obtained by training a model on the rest of the complete data. To reduce the imputation’s randomness, we will impute the missing variables 20 times for each MICE iteration instead of the usual practice of 10 times (Van Buuren & Groothuis-Oudshoorn, 2011). We tested the imputed and original distributions with the Kolmogorov-Smirnov test, which yielded a significance level of 0.5. Previously, Gibson et al. (2022) applied MICE to impute missing values of Active Galactic Nuclei. There, we demonstrated that this imputation method does not add any additional uncertainty to the variable’s original distribution and helps strengthen the final training set used for the statistical learning application by enlarging it.
@@ 3 Results
3.1 The Statistical learning results
We present the results derived with the ensemble method, the SuperLearner, which has the advantage of combining several statistical learning methods into a single model and leveraging the predictive power of each model. This is the first time SuperLearner has been applied to GRBs. SuperLearner computes coefficients for each learner based on their performance within the ensemble. Each model and the Superlearner are discussed in §A.1. In the left panel of Fig. 3, we show the observed redshift ($z_{obs}$) versus the predicted redshift ($z_{pred}$) via a 10-fold cross-validation technique averaged over 100 runs (see §A.3). The average mean squared error, MSE, is defined as the $<(z_{pred}-z_{obs})^{2}>$ where the symbol <>absent<>< > indicates the average between $z_{pred}$ vs $z_{obs}$. The obtained MSE is $0.21$, with a high Pearson correlation ($r$) of $0.93$ and a very small average bias, defined as the mean $<(z_{pred}-z_{obs})>$ of $0.02$.
@@ 3.1 The Statistical learning results
In the right panel of Fig. 3 we have shown the relative influence of the variables obtained with SuperLearner (see §A.3 for details). The relative importance of predictors is an average of the importance given by a local linear approximation of predictions. The variable $\log(NH)$ obtained the highest influence in predicting the redshift. This variable gives the NH obtained from an automated time-averaged spectral fit of the Swift XRT data (Evans et al., 2009). The high influence of this variable can be understood because the more distant a GRB is, the more neutral hydrogen it will encounter. Thus, the redshift of a GRB depends on the calculated $\log(NH)$ value. The second, third, fourth and fifth most predictive variables are $\log(F_{a})$, $\alpha$, $\log(Peakflux)$, $\log(T_{a})$. The variables related to prompt and the plateau emission, such as the time at the end of the plateau ($log(T_{a})$) and its correspondent flux $\log(F_{a})$, and the prompt $Peakflux$ are among the best predictors, thus recovering the key parameters of the three-dimensional Dainotti correlation (named the GRB fundamental plane relation), the correlation between the peak prompt luminosity, the luminosity at the end of the plateau emission and the rest end time of the plateau emission (Dainotti et al., 2016, 2017a, 2020a, 2020b, 2022).
@@ 3.2 Luminosity function and density rate evolution
To check whether our results can be used for deriving important astrophysical properties such as the GRB LF and the density rate evolution, we applied the Efron & Petrosian (Efron & Petrosian, 1992) method to derive these properties from both $z_{obs}$ and $z_{pred}$. To have more precise results, we selected 103 GRBs with $\delta z_{pred}/z_{pred}<1$. These GRBs are presented in the left panel of Fig. 3. We have analyzed the LF and density rate evolution using the method presented by Petrosian et al. (2015) and Dainotti et al. (2013c, 2015b). According to their prescription, the sample of 103 GRBs was further trimmed to 97 GRBs based on a flux limit of $1.25\times 10^{-14}$ erg cm-2 (see §A.5 and Fig.10).
@@
However, the values of these luminosity breaks are lower than those in Pescalli et al. (2016); Dong et al. (2022); Yu et al. (2015); Petrosian et al. (2015); Wanderman & Piran (2015, 2010), where the average log luminosity break is $51$. This is because we use plateau luminosity for deriving the LF. The difference between the $\log$ flux of the prompt and plateau is of the order of $5$. This estimate is taken as the difference between between an average of the logarithm of the peak flux of the prompt emission and the logarithm of flux at the end of the plateau emission in our sample. If we account for this in our luminosity break, then our values aligns with the literature.
The density rates (see Fig. 5) are calculated using the same 97 GRBs as in the LF. The density rate are computed using the following formula:
${\dot{\rho}}(z)=Z[d{\dot{\sigma}}(z)/dz]/[dV(z)/dz]$ (1)
where Z = z+1, $dV(z)/dz$ is the comoving volume element, and $\dot{\sigma}$(z) is the cumulative of the number of GRBs as a function of redshift. We consider the years of observations of these 97 GRBs from 24th January 2002 until 4th February 2021 (19.04 years). Since GRBs in our catalog majorly come from Swift (88%), then we also consider the field of view (FoV), and we remove the time in which Swift-BAT is not observing due to the South Atlantic Anomalies (SAA). The FoV is 17% of the whole sky, and the time of observations without the SAA for Swift-BAT is at 75.8% (Burns et al., 2016). Furthermore, our sample GRBs have optical plateaus and constitute a 36% fraction of the total sample. Thus, the density rate is adjusted to take this into account.
@@
The plots compare our density rate and those in the literature (upper panel of Fig. 5) and our density rate with SFR (lower panel of Fig. 5). We removed 7 GRBs which obtained a density rate going into the negative value within the error bars. This is due to the large errorbars of the parameters involved. Thus, the density rate plot features 90 GRBs as opposed to the 97 GRBs present in the luminosity function. We found the actual rate to be $\mathbf{8.47-9}$ Gpc $yr^{-1}$ for $2.31.9<z<$ 2.3 for the $z_{obs}$ and $z_{pred}$, respectively, without an arbitrary normalization. We have renormalized the rate (with a factor of 20) to account for our subset of LGRBs, compared to all LGRBs detected by Fermi and Swift, from the first detection of Fermi (14th July 2008) to 30th May 2023. We have quantitatively shown that the SFR does not match the GRB rate at low redshift ($z<1$). This discrepancy is of a factor of $\sim$3, comparing with Hopkins & Beacom (2006). The SFR that better fit our GRB data is Hopkins & Beacom (2006); Graham & Schady (2016), where the agreement runs from $z=0.8$ until $z=4.8$.
@@
The GRB rate does not follow the SFR at $z<1$, supporting the previous studies with a quantitative estimate. Although for most of the SFR compared here, the GRB rate does not follow the SFR at low-$z$, in the case of Thompson et al. (2006), there is no disagreement due to the very large errorbars. In general, the nature of this discrepancy can be due to selection biases (Petrosian et al., 2015) or to metallicity (Yoon et al., 2006; Ghirlanda & Salvaterra, 2022). The last two papers explain this excess due to the increased metallicity with cosmic time, which prevents a larger fraction of massive stars from producing GRBs. Ghirlanda & Salvaterra (2022) found that a metallicity threshold of $8.612+log(O/H)<=$ 8.6 accounts for the derived GRB efficiency evolution. They verified their results, computing the stellar masses of galaxies with metallicities below the threshold value found and compared them with the observed masses of a sample of GRB hosts claimed to be complete. The maximum and average stellar masses are consistent with the measured masses of the hosts and their redshift evolutions. However, these results depend on the claim of the completeness of the sample of the GRB hosts, which can be debated. Another relevant possibility is the low-$z$ GRBs have the same non-collapsar origin (Petrosian & Dainotti, 2023) of the SGRBs. The resulting LF and density rate evolutions are consistent with the ones in the literature and are also consistent between the ones derived from $z_{obs}$ and $z_{pred}$. However, due to the large uncertainties in the measurements both interpretations of Yoon et al. (2006) and Petrosian & Dainotti (2023) are viable. We have shown in the right panel of Fig. 4 that adding the estimated redshift improves the derivation of the break of the LF by 90%, namely $46.6\pm 0.3$ vs $47.2\pm 0.03$.
@@
Figure 5: The orange and purple dots show the density rates obtained from $z_{obs}$ and $z_{pred}$, respectively. Top Panel: the density rate evolution for our sample, compared with other published density rate evolutions for GRBs only. Bottom Panel: The density rate evolution for our GRB sample compared with the published SFR.